OpenStudy (anonymous):

What values for theta(0<= theat <= 2pi) satisfy the equation: 3sin theta = sin theta -1 a. pi/3, 5pi/3 b. 2pi/3, 4pi/3 c. pi/6, 3pi/6 7pi/6, 11pi/6 d. 7pi/6 ,11pi/6

geerky42 (geerky42):

\[3\sin\theta = \sin\theta - 1\]Correct?

geerky42 (geerky42):

Just isolate \(\sin\theta\)

OpenStudy (anonymous):

I worked it out to sinθ = -1/2. I'm honestly having a hard time grasping this stuff, so wondering what my next step would be?

geerky42 (geerky42):

Do you have unit circle you can use?

OpenStudy (anonymous):

Yes

geerky42 (geerky42):

OK, then find sine that equals -1/2 You should see what degrees you have as answer.

OpenStudy (anonymous):

240 degrees?

OpenStudy (anonymous):

I think I have it - is the answer B?

OpenStudy (anonymous):

thanks so much! If you're available, I have a lot of work to do tonight and would love help.

geerky42 (geerky42):



geerky42 (geerky42):



OpenStudy (anonymous):

I'm confused because there are a few different areas that have -1/2?

OpenStudy (anonymous):

What topic/area of Math is this? (Just so I'll know for future reference)